Please help.
Use the formula for compound interest. Solve for t when A=$10,800, P=$3600, and r=5%, and m=4

Answer by Nate(3500)

A = P(1 + r/m)^tm
10800 = 3600(1 + 0.05/4)^4t
3 = 1.0125^4t
log(3)/log(1.0125) = 4t
log(3)/(4*log(1.0125)) = t
About 22 Years and 1 Month